# A kernel conditional two-sample test

### v1

@@ D.3.2 Tensor factorization of Hilbert spaces and coefficients transformations
[…]
Definition D.10.
Let $\mathcal{H}$ and $\mathcal{G}$ be separable Hilbert spaces, with respective factorizations $\mathcal{H}\cong_{\iota_{\mathcal{H}}}H_{1}\otimes H_{2}$ and $\mathcal{G}\cong_{\iota_{\mathcal{G}}}G_{1}\otimes G_{2}$, and let $L\in\mathcal{L}_{\mathrm{b}}(\mathcal{H},\mathcal{G})$. We say that $L$ is $(\iota_{\mathcal{H}},\iota_{\mathcal{G}})$-tensor-factorizable [^6] For brevity, we refer to this property as simply “$(\iota_{\mathcal{H}},\iota_{\mathcal{G}})$-factorizability” in the remainder of the paper. if there exists a pair $(L_{1},L_{2})\in\mathcal{L}_{\mathrm{b}}(H_{1},G_{1})\times\mathcal{L}_{\mathrm{b}}(H_{2},G_{2})$ such that
@@
Figure 8: Action of a tensor-factorizable operator $L$ expressed as a commutative diagram. If $K$ is \@iaciUBD UBD kernel with block $K_{0}$, then the RKHS $\mathcal{H}_{K}$ and sampling operator $K(\cdot,x)^{\star}$ can be factorized as such with $L_{1}={K_{0}}(\cdot,x)^{\star}$ and $L_{2}=\mathrm{id}_{\mathcal{V}}$, with $\mathcal{V}:=H_{2}=G_{2}$, as per Proposition D.12.
Finally, we state the assumption the we require for the generalization of Theorem D.6 to hold.
Definition D.11.
Let $\mathcal{H}\cong_{\iota_{\mathcal{H}}}{\tilde{\mathcal{H}}}\otimes\mathcal{V}$ and $\mathcal{G}\cong_{\iota_{\mathcal{G}}}\tilde{\mathcal{G}}\otimes\mathcal{V}$ be separable Hilbert spaces with joint base space factorizations. Let $L\in\mathcal{L}_{\mathrm{b}}(\mathcal{H},\mathcal{G})$. We say that $L$ is an $(\iota_{\mathcal{H}},\iota_{\mathcal{G}})$-coefficients transformation if there exists $\tilde{L}\in\mathcal{L}_{\mathrm{b}}({\tilde{\mathcal{H}}},\tilde{\mathcal{G}})$ such that $(\tilde{L},\mathrm{id}_{\mathcal{V}})$ is an $(\iota_{\mathcal{H}},\iota_{\mathcal{G}})$-factorization of $L$.
@@
Proposition D.12.
Let $\mathcal{H}=\mathcal{H}_{K}$, where $K:\mathcal{X}\times\mathcal{X}\to\mathcal{L}_{\mathrm{b}}(\mathcal{G})$ is a kernel. Assume that $K$ is UBD with block decomposition ${K_{0}}:\mathcal{X}\times\mathcal{X}\to\mathcal{L}_{\mathrm{b}}(\tilde{\mathcal{G}})$ for the isometry $\iota_{\mathcal{G}}\in\mathcal{L}_{\mathrm{b}}(\tilde{\mathcal{G}}\otimes\mathcal{V},\mathcal{G})$, where $\tilde{\mathcal{G}}$ and $\mathcal{V}$ are separable Hilbert spaces. Then, there exists an isometry $\iota_{K}\in\mathcal{L}_{\mathrm{b}}(\mathcal{H}_{{K_{0}}}\otimes\mathcal{V},\mathcal{H}_{K})$ such that $\mathcal{H}\cong_{\iota_{K}}\mathcal{H}_{{K_{0}}}\otimes\mathcal{V}$, $\mathcal{G}\cong_{\iota_{\mathcal{G}}}\tilde{\mathcal{G}}\otimes\mathcal{V}$, and the family $L=(L_{x})_{x\in\mathcal{X}}=(K(\cdot,x)^{\star})_{x\in\mathcal{X}}$ is co-$(\iota_{K},\iota_{\mathcal{G}})$-factorizable and consists of coefficients transformations
@@
00 0 $=\left\langle\left({K_{0}}(\cdot,x)^{\star}\otimes\mathrm{id}_{\mathcal{V}}\right)u,g\otimes v_{p}\right\rangle_{\tilde{\mathcal{G}}\otimes\mathcal{V}}=\left\langle\sum_{n\in I_{K_{0}}}\alpha_{n,p}\cdot h_{n}(x),~{}g\right\rangle_{\tilde{\mathcal{G}}},$
@@
D.3.3 Self-normalized concentration of vector-valued martingales with coefficients-transforming measurements
The main result of this section is the following theorem, which is the generalized version of Theorem D.6. We name the relevant assumptions for future reference.
Assumption D.13.
Under the notations of Section D.1, $\mathcal{H}$ and $\mathcal{G}$ admit the factorizations with joint base space $\mathcal{H}\cong_{\iota_{\mathcal{H}}}{\tilde{\mathcal{H}}}\otimes\mathcal{V}$ and $\mathcal{G}\cong_{\iota_{\mathcal{G}}}\tilde{\mathcal{G}}\otimes\mathcal{V}$, $(L_{n})_{n\in{\mathbb{N}_{>0}}}$ is co-$(\iota_{\mathcal{H}},\iota_{\mathcal{G}})$-factorizable and consists of coefficient transformations, a.s., and $(\tilde{L}_{n},\mathrm{id}_{\mathcal{V}})$ is an $(\iota_{\mathcal{H}},\iota_{\mathcal{G}})$-factorization of $L_{n}$ with $\tilde{L}_{n}\tilde{L}_{n}^{\star}$ of trace class, a.s. and for all $n\in{\mathbb{N}_{>0}}$.
Theorem D.14.
Under the notations of Section D.1, assume Assumptions D.5 and D.13 hold. Assume that $\eta$ is $1$-subgaussian conditionally on $\mathcal{F}$. Let $Q\in\mathcal{L}_{\mathrm{b}}(\mathcal{V})$ be self-adjoint, positive semi-definite, and trace-class, and $\tilde{V}\in\mathcal{L}({\tilde{\mathcal{H}}})$ be self-adjoint, positive definite, diagonal, and with bounded inverse. Then, $\tilde{M}_{N}\tilde{V}^{-1}\tilde{M}_{N}^{\star}$ is a.s. of trace class for all $N\in{\mathbb{N}_{>0}}$ and, for all $\delta\in(0,1)$,
$\mathbb{P}\left[\forall N\in{\mathbb{N}_{>0}},\lVert\iota_{\mathcal{H}}^{-1}S_{N}\rVert_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q}\leq\sqrt{2\mathrm{Tr}(Q)\ln\left(\frac{1}{\delta}\left[\det(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\tilde{M}_{N}\tilde{V}^{-1}\tilde{M}_{N}^{\star})\right]^{1/2}\right)}\right]\geq 1-\delta,$ (26)
@@
$\begin{split}\tilde{M}_{N}:\tilde{h}\in{\tilde{\mathcal{H}}}\mapsto(\tilde{L}_{n}^{\star}\tilde{h})_{n\in\left[N\right]}\in\tilde{\mathcal{G}}^{N},\quad\text{and}\quad\tilde{V}_{N}=\tilde{M}_{N}^{\star}\tilde{M}_{N}\in\mathcal{L}_{\mathrm{b}}({\tilde{\mathcal{H}}}).\end{split}$ (27)
We prove this theorem first under the assumption that $\tilde{V}^{-1}$ is of trace class rather than $\tilde{L}_{n}\tilde{L}_{n}^{\star}$ is, as it enables detailing all of the arguments in a technically simpler setup. We then generalize the proof to handle the setup of Theorem D.14.
Lemma D.15.
Under the same setup and assumptions as Theorem D.14, do not assume that $\tilde{L}_{n}\tilde{L}_{n}^{\star}$ is of trace class for all $n\in{\mathbb{N}_{>0}}$, but rather that $\tilde{V}^{-1}$ is of trace class. Then, for all $\delta\in(0,1)$, (26) holds.
To show this, we follow the method announced in Section D.3.1.
Lemma D.16.
Under the setup and assumptions as Lemma D.15, define for all $N\in{\mathbb{N}_{>0}}$
@@
~$\int_{{\tilde{\mathcal{H}}}}G_{N}[\tilde{h}\otimes v]\mathrm{d}\mathcal{N}(\tilde{h}\mid 0,\tilde{V}^{-1})=\left[\det\left(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\lVert v\rVert^{2}\tilde{M}_{N}\tilde{V}^{-1}\tilde{M}_{N}\right)\right]^{-1/2}\exp\left[\frac{1}{2}\lVert\tilde{S}_{N}(v)\rVert^{2}_{(\tilde{V}+\lVert v\rVert^{2}\tilde{V}_{N})^{-1}}\right],$ (29)
@@
where $\mathrm{Cont}(v)\in\mathcal{L}_{\mathrm{b}}({\tilde{\mathcal{H}}}\otimes\mathcal{V},{\tilde{\mathcal{H}}})$ is the tensor contraction of $v$, that is, the unique bounded operator that satisfies $\mathrm{Cont}(v)(u\otimes w)=\langle v,w\rangle\cdot u$ for all $u\otimes w\in{\tilde{\mathcal{H}}}\otimes\mathcal{V}$.
@@
~$\int_{{\tilde{\mathcal{H}}}}G_{N}(\tilde{h}\otimes v)\mathrm{d}\mathcal{N}(\tilde{h}\mid 0,\tilde{V}^{-1})$ =~$=\int_{{\tilde{\mathcal{H}}}}\exp\left[\langle\tilde{h},\tilde{S}_{N}(v)\rangle_{\tilde{\mathcal{H}}}-\frac{1}{2}\lVert v\rVert^{2}_{\mathcal{V}}\lVert\tilde{h}\rVert^{2}_{\tilde{V}_{N}}\right]\mathrm{d}\mathcal{N}(\tilde{h}\mid 0,\tilde{V}^{-1})$
=~$=\int_{{\tilde{\mathcal{H}}}}\exp\left[\langle u,\tilde{S}_{N}(\bar{v})\rangle_{\tilde{\mathcal{H}}}-\frac{1}{2}\lVert u\rVert^{2}_{\tilde{V}_{N}}\right]\mathrm{d}\mathcal{N}(u\mid 0,\lVert v\rVert_{\mathcal{V}}\tilde{V}^{-1}),$
@@
$\left[\det\left(\mathrm{id}_{\tilde{\mathcal{H}}}+\lVert v\rVert_{\mathcal{V}}^{2}\tilde{V}^{-1/2}\tilde{V}_{N}\tilde{V}^{-1/2}\right)\right]^{-1/2}\exp\left[\frac{1}{2}\left\lVert\left(\mathrm{id}_{\tilde{\mathcal{H}}}+\lVert v\rVert_{\mathcal{V}}^{2}\tilde{V}^{-1/2}\tilde{V}_{N}\tilde{V}^{-1/2}\right)^{-1/2}\lVert v\rVert_{\mathcal{V}}\tilde{V}^{-1/2}\tilde{S}_{N}(\bar{v})\right\rVert^{2}_{{\tilde{\mathcal{H}}}}\right].$
@@
Lemma D.18.
Under the setup and assumptions of Lemma D.15, let $G_{N}$ be defined for all $N\in{\mathbb{N}_{>0}}$ as in (28). Assume that $Q\neq 0$, and let $\mu_{Q}$ be a probability measure on $\mathcal{V}$ as given in Lemma D.17. Then,
$\int_{\mathcal{V}}\int_{{\tilde{\mathcal{H}}}}G_{N}(\tilde{h}\otimes v)\mathrm{d}\mathcal{N}(\tilde{h}\mid 0,\tilde{V}^{-1})\mathrm{d}\mu_{Q}(v)\geq\left[\det\left(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\tilde{M}_{N}\tilde{V}^{-1}\tilde{M}_{N}^{\star}\right)\right]^{-1/2}\exp\left[\frac{1}{2\mathrm{Tr}(Q)}\lVert\iota_{\mathcal{H}}^{-1}S_{N}\rVert^{2}_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q}\right].$
@@
It follows from Lemma D.16 and the fact that $\mu_{Q}$ has support in $\mathbb{S}_{\mathcal{V}}^{1}$ that
@@
$=\int_{\mathbb{S}^{1}_{\mathcal{V}}}\sum_{n,m=1}^{N}\sum_{i,j\in J}\langle(\tilde{V}+\tilde{V}_{N})^{-1}\mathrm{Cont}(v)(\tilde{L}_{n}^{\star}\otimes\mathrm{id}_{\mathcal{V}})(e_{i}^{(n)}\otimes f_{i}^{(n)}),\mathrm{Cont}(v)(\tilde{L}_{m}^{\star}\otimes\mathrm{id}_{\mathcal{V}})(e_{j}^{(m)}\otimes f_{j}^{(m)})\rangle_{{\tilde{\mathcal{H}}}}\mathrm{d}\mu_{Q}(v)$
@@
$=\frac{1}{\mathrm{Tr}(Q)}\sum_{n,m=1}^{N}\sum_{i,j\in J}\left\langle\left[(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q\right](\tilde{L}_{n}^{\star}\otimes\mathrm{id}_{\mathcal{V}})(e_{i}^{(n)}\otimes f_{i}^{(n)}),(\tilde{L}_{m}^{\star}\otimes\mathrm{id}_{\mathcal{V}})(e_{j}^{(m)}\otimes f_{j}^{(m)})\right\rangle_{{\tilde{\mathcal{H}}}\otimes\mathcal{V}}$
@@
$\int_{\mathbb{S}^{1}_{\mathcal{V}}}\lVert\tilde{S}_{N}(v)\rVert^{2}_{(\tilde{V}+\tilde{V}_{N})^{-1}}\mathrm{d}\mu_{Q}(v)=\frac{1}{\mathrm{Tr}(Q)}\left\langle\left[(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q\right]\iota_{\mathcal{H}}^{-1}S_{N},\iota_{\mathcal{H}}^{-1}S_{N}\right\rangle_{{\tilde{\mathcal{H}}}\otimes\mathcal{V}}=\frac{1}{\mathrm{Tr}(Q)}\lVert\iota_{\mathcal{H}}^{-1}S_{N}\rVert_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q},$
@@
Lemma D.19.
Under the setup and assumption of Lemma D.15, define $G_{N}$ as in (28). Then, for all $h\in{\tilde{\mathcal{H}}}\otimes\mathcal{V}$, $(G_{N}(h))_{N}$ is an $\mathcal{F}$-supermartingale and $\mathbb{E}[G_{N}(h)]\leq 1$ for all $N\in{\mathbb{N}_{>0}}$.
@@
$\mathbb{E}[G_{N}(h)\mid\mathcal{F}_{N-1}]$ $=\mathbb{E}\left[\exp\left(\sum_{n=1}^{N}\langle\iota_{\mathcal{H}}h,L_{n}^{\star}\eta_{n}\rangle_{\mathcal{H}}-\frac{1}{2}\lVert\iota_{\mathcal{H}}h\rVert_{L_{n}^{\star}L_{n}}\right)\middle|\mathcal{F}_{N-1}\right]$
$=\mathbb{E}\left[\exp\left(-\frac{1}{2}\langle\iota_{\mathcal{H}}h,L_{N}^{\star}L_{N}\iota_{\mathcal{H}}h\rangle_{\mathcal{H}}\right)\exp(\langle L_{N}\iota_{\mathcal{H}}h,\eta_{N}\rangle_{\mathcal{G}})G_{N-1}(h)~{}\middle|~{}\mathcal{F}_{N-1}\right]$
$=\exp\left(-\frac{1}{2}\lVert L_{N}\iota_{\mathcal{H}}h\rVert^{2}_{\mathcal{G}}\right)\mathbb{E}\left[\exp(\langle L_{N}\iota_{\mathcal{H}}h,\eta_{N}\rangle_{\mathcal{G}})~{}\middle|~{}\mathcal{F}_{N-1}\right]\cdot G_{N-1}(h)$
$\leq\exp\left(-\frac{1}{2}\lVert L_{N}\iota_{\mathcal{H}}h\rVert^{2}_{\mathcal{G}}\right)\exp\left(\frac{1}{2}\lVert L_{N}\iota_{\mathcal{H}}h\rVert^{2}_{\mathcal{G}}\right)\cdot G_{N-1}(h)$
$=G_{N-1}(h),$
@@
Proof of Lemma D.15.
First, the claim is trivial if $Q=0$. Assume that $Q\neq 0$. Under the assumptions of the lemma, we apply Lemma D.19 and obtain that $\mathbb{E}[G_{N}(h)]\leq 1$ for all $N\in{\mathbb{N}_{>0}}$ and $h\in{\tilde{\mathcal{H}}}\otimes\mathcal{V}$. In particular, this is the case for any $h$ of the form $h=\tilde{h}\otimes v$, where $\tilde{h}\in{\tilde{\mathcal{H}}}$ and $v\in\mathcal{V}$. Therefore, by Fubini’s theorem, the process $(G_{N})_{N}$ defined for all $N\in{\mathbb{N}_{>0}}$ as
@@
Lemma D.18 now shows that
$\mathbb{P}\left[\forall N\in{\mathbb{N}_{>0}},\left(\det\left(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\tilde{M}_{N}\tilde{V}^{-1}M_{N}\right)\right)^{-1/2}\exp\left(\frac{1}{2\mathrm{Tr}(Q)}\lVert\iota_{\mathcal{H}}^{-1}S_{N}\rVert^{2}_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q}\right)\leq\frac{1}{\delta}\right]\geq 1-\delta.$

### v2

@@ E.3.2 Tensor factorization of Hilbert spaces and coefficients transformations
The existence of an isometry $\iota_{\mathcal{H}}$ such that $\mathcal{H}\cong_{\iota_{\mathcal{H}}}H_{1}\otimes H_{2}$ for given (separable) $\mathcal{H}$, $H_{1}$, and $H_{2}$ can be assessed based on purely dimensional arguments: there exists such an $\iota_{\mathcal{H}}$ if, and only if, $\dim(\mathcal{H})=\dim(H_{1})\cdot\dim(H_{2})$ (with $0\cdot\infty:=0$ and $\infty\cdot\infty:=\infty$). For instance, there always exists the trivial factorization of a Hilbert space $\mathcal{H}\cong_{\iota_{1}}\mathbb{R}\otimes\mathcal{H}$, and if $n=p\cdot q$, $\mathbb{R}^{n}\cong_{\iota_{2}}\mathbb{R}^{p}\otimes\mathbb{R}^{q}$. It follows that the relationship $\mathcal{H}\cong_{\iota_{\mathcal{H}}}H_{1}\otimes H_{2}$ does not impose a lot of constraints on any of the spaces involved. The main interest of the notion is when relating it to operators, as operators between Hilbert spaces may have a “nice” form when seen as operators between their factorizations, as illustrated in Figure 8:
Definition E.10.
Let $\mathcal{H}$ and $\mathcal{G}$ be separable Hilbert spaces, with respective factorizations $\mathcal{H}\cong_{\iota_{\mathcal{H}}}H_{1}\otimes H_{2}$ and $\mathcal{G}\cong_{\iota_{\mathcal{G}}}G_{1}\otimes G_{2}$, and let $L\in\mathcal{L}_{\mathrm{b}}(\mathcal{H},\mathcal{G})$. We say that $L$ is $(\iota_{\mathcal{H}},\iota_{\mathcal{G}})$-tensor-factorizable [^5] For brevity, we refer to this property as simply $(\iota_{\mathcal{H}},\iota_{\mathcal{G}})$-factorizability” in the remainder of the paper. if there exists a pair $(L_{1},L_{2})\in\mathcal{L}_{\mathrm{b}}(H_{1},G_{1})\times\mathcal{L}_{\mathrm{b}}(H_{2},G_{2})$ such that
@@
Figure 8: Action of a tensor-factorizable operator $L$ expressed as a commutative diagram. If $K$ is \@iaciUBD UBD kernel with block $K_{0}$, then the RKHS $\mathcal{H}_{K}$ and sampling operator $K(\cdot,x)^{\star}$ can be factorized as such with $L_{1}={K_{0}}(\cdot,x)^{\star}$ and $L_{2}=\mathrm{id}_{\mathcal{V}}$, with $\mathcal{V}:=H_{2}=G_{2}$, as per Proposition E.12.
Finally, we state the assumption the we require for the generalization of Theorem E.6 to hold.
Definition E.11.
[…]
@@
Proposition E.12.
[…]
@@
$0$ $=\left\langle\left({K_{0}}(\cdot,x)^{\star}\otimes\mathrm{id}_{\mathcal{V}}\right)u,g\otimes v_{p}\right\rangle_{\tilde{\mathcal{G}}\otimes\mathcal{V}}=\left\langle\sum_{n\in I_{K_{0}}}\alpha_{n,p}\cdot h_{n}(x),~g\right\rangle_{\tilde{\mathcal{G}}}$,
@@
E.3.3 Self-normalized concentration of vector-valued martingales with coefficients-transforming measurements
The main result of this section is the following theorem, which is the generalized version of Theorem E.6. We name the relevant assumptions for future reference.
Assumption E.13.
Under the notations of Section E.1, $\mathcal{H}$ and $\mathcal{G}$ admit the factorizations with joint base space $\mathcal{H}\cong_{\iota_{\mathcal{H}}}{\tilde{\mathcal{H}}}\otimes\mathcal{V}$ and $\mathcal{G}\cong_{\iota_{\mathcal{G}}}\tilde{\mathcal{G}}\otimes\mathcal{V}$, $(L_{n})_{n\in{\mathbb{N}_{>0}}}$ is co-$(\iota_{\mathcal{H}},\iota_{\mathcal{G}})$-factorizable and consists of coefficient transformations, a.s., and $(\tilde{L}_{n},\mathrm{id}_{\mathcal{V}})$ is an $(\iota_{\mathcal{H}},\iota_{\mathcal{G}})$-factorization of $L_{n}$ with $\tilde{L}_{n}\tilde{L}_{n}^{\star}$ of trace class, a.s. and for all $n\in{\mathbb{N}_{>0}}$.
Theorem E.14.
[…]
$\forall N\in{\mathbb{N}_{>0}},\lVert\iota_{\mathcal{H}}^{-1}S_{N}\rVert_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q}\leq\sqrt{2\mathrm{Tr}(Q)\ln\left(\frac{1}{\delta}\left[\det(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\tilde{M}_{N}\tilde{V}^{-1}\tilde{M}_{N}^{\star})\right]^{1/2}\right)},$ (26)
@@
[…]
We prove this theorem first under the assumption that $\tilde{V}^{-1}$ is of trace class rather than $\tilde{L}_{n}\tilde{L}_{n}^{\star}$ is, as it enables detailing all of the arguments in a technically simpler setup. We then generalize the proof to handle the setup of Theorem E.14.
Lemma E.15.
Under the same setup and assumptions as Theorem E.14, do not assume that $\tilde{L}_{n}\tilde{L}_{n}^{\star}$ is of trace class for all $n\in{\mathbb{N}_{>0}}$, but rather that $\tilde{V}^{-1}$ is of trace class. Then, for all $\delta\in(0,1)$, (26) holds.
To show this, we follow the method announced in Section E.3.1.
Lemma E.16.
Under the setup and assumptions as Lemma E.15, define for all $N\in{\mathbb{N}_{>0}}$
@@
$\begin{split}&\int_{{\tilde{\mathcal{H}}}}G_{N}[\tilde{h}\otimes v]\mathrm{d}\mathcal{N}(\tilde{h}\mid 0,\tilde{V}^{-1})\\ &=\left[\det\left(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\lVert v\rVert^{2}\tilde{M}_{N}\tilde{V}^{-1}\tilde{M}_{N}\right)\right]^{-1/2}\exp\left[\frac{1}{2}\lVert\tilde{S}_{N}(v)\rVert^{2}_{(\tilde{V}+\lVert v\rVert^{2}\tilde{V}_{N})^{-1}}\right],\end{split}$ (29)
@@
[…]
@@
$\int_{{\tilde{\mathcal{H}}}}G_{N}(\tilde{h}\otimes v)\mathrm{d}\mathcal{N}(\tilde{h}\mid 0,\tilde{V}^{-1})$ $=\int_{{\tilde{\mathcal{H}}}}\exp\left[\langle\tilde{h},\tilde{S}_{N}(v)\rangle_{\tilde{\mathcal{H}}}-\frac{1}{2}\lVert v\rVert^{2}_{\mathcal{V}}\lVert\tilde{h}\rVert^{2}_{\tilde{V}_{N}}\right]\mathrm{d}\mathcal{N}(\tilde{h}\mid 0,\tilde{V}^{-1})$
$=\int_{{\tilde{\mathcal{H}}}}\exp\left[\langle u,\tilde{S}_{N}(\bar{v})\rangle_{\tilde{\mathcal{H}}}-\frac{1}{2}\lVert u\rVert^{2}_{\tilde{V}_{N}}\right]\mathrm{d}\mathcal{N}(u\mid 0,\lVert v\rVert_{\mathcal{V}}\tilde{V}^{-1}),$
@@
$\left[\det\left(\mathrm{id}_{\tilde{\mathcal{H}}}+\lVert v\rVert_{\mathcal{V}}^{2}\tilde{V}^{-1/2}\tilde{V}_{N}\tilde{V}^{-1/2}\right)\right]^{-1/2}$
$\times\exp\left[\frac{1}{2}\left\lVert\left(\mathrm{id}_{\tilde{\mathcal{H}}}+\lVert v\rVert_{\mathcal{V}}^{2}\tilde{V}^{-1/2}\tilde{V}_{N}\tilde{V}^{-1/2}\right)^{-1/2}\lVert v\rVert_{\mathcal{V}}\tilde{V}^{-1/2}\tilde{S}_{N}(\bar{v})\right\rVert^{2}_{{\tilde{\mathcal{H}}}}\right].$
@@
Lemma E.18.
Under the setup and assumptions of Lemma E.15, let $G_{N}$ be defined for all $N\in{\mathbb{N}_{>0}}$ as in (28). Assume that $Q\neq 0$, and let $\mu_{Q}$ be a probability measure on $\mathcal{V}$ as given in Lemma E.17. Then,
$\geq\left[\det\left(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\tilde{M}_{N}\tilde{V}^{-1}\tilde{M}_{N}^{\star}\right)\right]^{-1/2}\exp\left[\frac{1}{2\mathrm{Tr}(Q)}\lVert\iota_{\mathcal{H}}^{-1}S_{N}\rVert^{2}_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q}\right].$
@@
It follows from Lemma E.16 and the fact that $\mu_{Q}$ has support in $\mathbb{S}_{\mathcal{V}}^{1}$ that
@@
$=\int_{\mathbb{S}^{1}_{\mathcal{V}}}\sum_{n,m=1}^{N}\sum_{i,j\in J}\langle(\tilde{V}+\tilde{V}_{N})^{-1}\mathrm{Cont}(v)(\tilde{L}_{n}^{\star}\otimes\mathrm{id}_{\mathcal{V}})(e_{i}^{(n)}\otimes f_{i}^{(n)}),$
$\qquad\qquad\mathrm{Cont}(v)(\tilde{L}_{m}^{\star}\otimes\mathrm{id}_{\mathcal{V}})(e_{j}^{(m)}\otimes f_{j}^{(m)})\rangle_{{\tilde{\mathcal{H}}}}\mathrm{d}\mu_{Q}(v)$
@@
$=\frac{1}{\mathrm{Tr}(Q)}\sum_{n,m=1}^{N}\sum_{i,j\in J}\left\langle\left[(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q\right](\tilde{L}_{n}^{\star}\otimes\mathrm{id}_{\mathcal{V}})(e_{i}^{(n)}\otimes f_{i}^{(n)})\right$.,
$\left.\qquad\qquad(\tilde{L}_{m}^{\star}\otimes\mathrm{id}_{\mathcal{V}})(e_{j}^{(m)}\otimes f_{j}^{(m)})\right\rangle_{{\tilde{\mathcal{H}}}\otimes\mathcal{V}}$
@@
$\int_{\mathbb{S}^{1}_{\mathcal{V}}}\lVert\tilde{S}_{N}(v)\rVert^{2}_{(\tilde{V}+\tilde{V}_{N})^{-1}}\mathrm{d}\mu_{Q}(v)$ $=\frac{1}{\mathrm{Tr}(Q)}\left\langle\left[(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q\right]\iota_{\mathcal{H}}^{-1}S_{N},\iota_{\mathcal{H}}^{-1}S_{N}\right\rangle_{{\tilde{\mathcal{H}}}\otimes\mathcal{V}}$
$=\frac{1}{\mathrm{Tr}(Q)}\lVert\iota_{\mathcal{H}}^{-1}S_{N}\rVert_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q},$
@@
Lemma E.19.
Under the setup and assumption of Lemma E.15, define $G_{N}$ as in (28). Then, for all $h\in{\tilde{\mathcal{H}}}\otimes\mathcal{V}$, $(G_{N}(h))_{N}$ is an $\mathcal{F}$-supermartingale and $\mathbb{E}[G_{N}(h)]\leq 1$ for all $N\in{\mathbb{N}_{>0}}$.
@@
$\mathbb{E}[G_{N}(h)\mid\mathcal{F}_{N-1}]$
$\quad=\mathbb{E}\left[\exp\left(\sum_{n=1}^{N}\langle\iota_{\mathcal{H}}h,L_{n}^{\star}\eta_{n}\rangle_{\mathcal{H}}-\frac{1}{2}\lVert\iota_{\mathcal{H}}h\rVert_{L_{n}^{\star}L_{n}}\right)\middle|\mathcal{F}_{N-1}\right]$
$\quad=\mathbb{E}\left[\exp\left(-\frac{1}{2}\langle\iota_{\mathcal{H}}h,L_{N}^{\star}L_{N}\iota_{\mathcal{H}}h\rangle_{\mathcal{H}}\right)\exp(\langle L_{N}\iota_{\mathcal{H}}h,\eta_{N}\rangle_{\mathcal{G}})G_{N-1}(h)~\middle|~\mathcal{F}_{N-1}\right]$
$\quad=\exp\left(-\frac{1}{2}\lVert L_{N}\iota_{\mathcal{H}}h\rVert^{2}_{\mathcal{G}}\right)\mathbb{E}\left[\exp(\langle L_{N}\iota_{\mathcal{H}}h,\eta_{N}\rangle_{\mathcal{G}})~\middle|~\mathcal{F}_{N-1}\right]\cdot G_{N-1}(h)$
$\quad\leq\exp\left(-\frac{1}{2}\lVert L_{N}\iota_{\mathcal{H}}h\rVert^{2}_{\mathcal{G}}\right)\exp\left(\frac{1}{2}\lVert L_{N}\iota_{\mathcal{H}}h\rVert^{2}_{\mathcal{G}}\right)\cdot G_{N-1}(h)$
$\quad=G_{N-1}(h),$
@@
Proof of Lemma E.15.
First, the claim is trivial if $Q=0$. Assume that $Q\neq 0$. Under the assumptions of the lemma, we apply Lemma E.19 and obtain that $\mathbb{E}[G_{N}(h)]\leq 1$ for all $N\in{\mathbb{N}_{>0}}$ and $h\in{\tilde{\mathcal{H}}}\otimes\mathcal{V}$. In particular, this is the case for any $h$ of the form $h=\tilde{h}\otimes v$, where $\tilde{h}\in{\tilde{\mathcal{H}}}$ and $v\in\mathcal{V}$. Therefore, by Fubini’s theorem, the process $(G_{N})_{N}$ defined for all $N\in{\mathbb{N}_{>0}}$ as
@@
Lemma E.18 now shows that it holds with probability at least $1-\delta$ that
$\forall N\in{\mathbb{N}_{>0}},\left(\det\left(\mathrm{id}_{\tilde{\mathcal{G}}^{N}}+\tilde{M}_{N}\tilde{V}^{-1}M_{N}\right)\right)^{-1/2}\exp\left(\frac{1}{2\mathrm{Tr}(Q)}\lVert\iota_{\mathcal{H}}^{-1}S_{N}\rVert^{2}_{(\tilde{V}+\tilde{V}_{N})^{-1}\otimes Q}\right)\leq\frac{1}{\delta}.$